# A Statistician’s Overview of Mechanistic-Informed Modeling

### v1

@@ 6 Simulation Experiment: GMID-DSTM
where relatively informative priors are assigned for $\sigma^{2}_{d}$, $\sigma^{2}_{\nu}$, and $\ell_{\nu}$ with $\mu_{d}=\mu_{\nu}=\mu_{\ell}=0,\gamma_{d}=\gamma_{\nu}=\sigma^{2}_{\ell}=1$, and truncation bounds $\ell_{d}=0.1,u_{d}=0.3,\ell_{\nu}=0.02,u_{\nu}=0.1,\ell_{\ell}=0.05,$ and $u_{\ell}=0.2$. This is to address the potential identifiability issue between the measurement error $\sigma^{2}_{d}$ and the discrepancy scale parameter, as well as the ability to recover Gaussian process covariance parameters in a fixed domain setting (an issue which has been frequently noted in the spatial statistics literature as for instance in Gelfand et al., (2010)). In contrast, we assign relatively non-informative priors to $\bm{\beta}$, $\lambda$, the neural network weight and bias parameters $\bm{\theta}_{W}$, with $\bm{\mu}_{\bm{\beta}}=\bm{\mu}_{W}=\bm{0}$, $c_{W}=\sigma^{2}_{\lambda}=1$, $\mu_{\lambda}=0$, and $c_{\bm{\beta}}=10$. The neural network for ${\bm{u}}_{NN,t}$ consists of $L=3$ hidden layers, each with 16 hidden units.
Figure 3: Violin plots of the BHM model parameter posterior distributions, with the true value indicated by a dashed red line. From left to right, the plots correspond to $\sigma^{2}_{\nu}$, $\ell_{\nu}$, $\lambda$, $\sigma^{2}_{d}$, $\beta_{1}$, and $\beta_{2}$.
Figure 4: Left: true $\tilde{\bm{u}}_{t}$; Second from the left: Predicted $\tilde{\bm{u}}_{t}$ from $u_{NN}(s,t)$; Second from the right: Posterior means (dashed line) and corresponding 95% credible intervals (shaded region) for predicted values of a marginal spatial slice of $\tilde{\bm{u}}_{t}$ at times $t=0$, $2.5$, and $5$ compared to the true process (solid line); Right: Posterior means (dashed line) and corresponding 95% credible intervals (shaded region) for a marginal time slice of the predicted values of $\tilde{\bm{u}}_{t}$ at spatial locations $s=-3.14$, $-1.13$, and $1.01$ compared to the true process (solid line). In the right two plots, the dots correspond to spatial locations or time points when data are observed.
Figure 5: First Row, First Column: True values of the spatio-temporal discrepancy term $\bm{\nu}_{t}$; Second Row, First Column: Corresponding posterior prediction of $\bm{\nu}_{t}$; First Row, Columns 2-4: Spatial slice of posterior mean (dashed line) and 95% credible intervals (shaded region) for times $t=0$, $2$, and $t=5$ compared to the true process (solid line); Second Row, Columns 2-4: Time slice of posterior mean (dashed line) and 95% credible intervals (shaded region) for spatial locations $s=-3$, $-1$, and 1 compared to the true process (solid line). For all plots in columns 2-4, the dots correspond to spatial locations or time points when data is observed.
@@
As discussed in Section 4.2, the specification of the B-PINN model in this BHM requires PDE residuals, boundary condition data, and initial condition data. Specifically, we can write the following conditional data distributions that have the effect of updating the prior distributions for weights $\bm{\theta}_{W}$.
[…]
$[D_{bc+}|\sigma^{2}_{bc},\bm{\theta}_{W}]$ $=\text{Gau}(0;u_{NN}(\pi,t;\bm{\theta}_{W}),\sigma^{2}_{bc}),\quad(\text{Right boundary condition})$
$[D_{bc-}|\sigma^{2}_{bc},\bm{\theta}_{W}]$ $=\text{Gau}(0;u_{NN}(-\pi,t;\bm{\theta}_{W}),\sigma^{2}_{bc}),\quad(\text{Left boundary condition})$
$[D_{ic}|\sigma^{2}_{ic},\bm{\theta}_{W}]$ $=\text{Gau}(\exp(-s^{2});u_{NN}(s,0;\bm{\theta}_{W}),\sigma^{2}_{ic}),\quad(\text{Initial condition}),$
@@
where $\sigma^{2}_{r}$, $\sigma^{2}_{bc}$, and $\sigma^{2}_{ic}$ govern the extent to which the neural network $u_{NN}$ is constrained to satisfy Burgers’ equation along with the initial conditions and boundary conditions. Assuming reasonably strong prior information for the PDE, boundary conditions, and initial conditions, we set $\sigma^{2}_{r}=0.05^{2}$ and $\sigma^{2}_{bc}=\sigma^{2}_{ic}=0.01^{2}$. We emphasize that assuming the residual data are zero (e.g., $D_{r}=0$) further constrains the prior on $\bm{\theta}_{W}$ and these residuals are not actual measurements of the process.
@@
After specifying the prior distributions and the neural network architecture, we draw 2,000 posterior samples using HMC with the No-U-Turn Sampler (Hoffman & Gelman,, 2014), discarding the first 1,000 as burn-in and using a tree depth of 10. This is implemented using the “NumPyro" (Phan et al.,, 2019) and “JAX" (Bradbury et al.,, 2018) packages in Python and takes approximately 90 minutes on an Apple MacBook M1 Pro with 16GB RAM.
@@ 6.1 Results
Figures 4 and 5 compare the true values and predictions of ${\bm{u}}_{d,t}$ and $\bm{\nu}_{t}$, respectively. As shown in Figure 4, the predicted ${\bm{u}}_{d,t}$ space-time field and spatial and temporal marginal plots align well with the true values, despite the substantial amount of missing data. This is one of the primary advantages of incorporating mechanistic information into the model. In contrast, Figure 5 demonstrates that accurately predicting the discrepancy term, $\bm{\nu}_{t}$, is more challenging and subject to greater uncertainty, although the posterior mean (dashed line in the spatial and temporal slice plots) still reasonably matches the true process (solid line). Combining posterior samples from each term, Figure 6 shows the posterior mean of the overall process $\bm{u}_{t}$. While there are slight deviations from the true $\bm{u}_{t}$, the posterior mean successfully captures the underlying spatio-temporal pattern, demonstrating the flexibility of the model.
@@ 7 Discussion
The incorporation of mechanistic information in data-driven models has been of interest to scientists for decades. There are numerous paradigms for such data-driven approaches, ranging from variational data assimilation, online filtering, physical-statistical models and probabilistic numerical methods. Recently, there has been a resurgence of interest in these approaches because of the success of using deep neural networks with physical constraints. Here, we review these different approaches and show some connections, followed by presenting a general BHM framework that can accommodate each of these paradigms. We illustrate this with a simulation of a process that is informed by the classic Burgers’ equation nonlinear PDE, and demonstrate we can recover important components of the model (fixed effect mean parameters, the Burgers’ diffusion parameter, and a discrepancy process) given uncertain and missing data.
@@
One novel contribution of this work has been the introduction of a mechanistically-informed BHM (GMID-DSTM) that considers a B-PINN for Burgers’ equation (Section 6). We expect that the same sort of approach can generalize to different processes motivated by alternate PDEs. In other cases, it may be preferred to embed a dynamical process that has spatio-temporal covariance, in contrast to when dynamics are encoded with a neural network. For example in Wikle et al., (2001), 2-D spatio-temporal wind velocity follows a vector-autoregressive process that is motivated by an underlying set of (linear) PDEs. However, for general nonlinear PDEs there may not be an autoregressive or general quadratic nonlinear (Wikle & Hooten,, 2010) representation or approximation of the latent dynamics. In such cases a mechanistic-BHM using a PINN approach may be preferred since automatic differentiation can be used for general, nonlinear PDEs. This could also be an advantage of using a model such as the one in Section 6 over PNM methods that lack a neural network representation, in which case it is not obvious that automatic differentiation would apply.
@@
From our exploration of these methods, in particular, the incorporation of a B-PINN within a more general BHM, there are several challenges with this approach that suggest additional research. Of primary concern is the challenge of computational efficiency when scaling this to a realistic environmental application. Although automatic differentiation is quite helpful for implementing PDE models, it can be slow, and there are well-known challenges to Bayesian learning of deep neural network parameters. In addition, real-world processes are often described by more complex dynamics than can be parameterized by a PDE (all models are wrong), and modeling the discrepancy becomes crucial. Our simulation studies have shown there can be significant identifiability issues with a flexible discrepancy process and the underlying mechanistic process without rough a priori guidance as to at least the space and time scales corresponding to each. It is also the case the data volume and orientation can make a difference in accurately estimating the various model components. Lastly, as with all deep neural implementations, there is typically no a priori guidance to inform the model architecture. These are very active areas of research and we are optimistic that these challenges can be met and these methods can be scaled to many realistic spatio-temporal problems.

### v2

@@ 6 Simulation Experiment: BHM-PINN
where relatively informative priors are assigned for $\sigma^{2}_{\delta}$ and $\ell_{\delta}$ with $\mu_{\delta}=\mu_{\ell}=0,\gamma^{2}_{\delta}=\sigma^{2}_{\ell}=1$, and truncation bounds $\ell_{\delta}=0.02,u_{\delta}=0.1,\ell_{\ell}=0.05,$ and $u_{\ell}=0.2$. These informative priors on the discrepancy parameters address the potential identifiability issue between the NN and discrepancy as well as the ability to recover Gaussian process covariance parameters in a fixed domain setting (an issue which has been frequently noted in the spatial statistics literature, e.g., gelfand2010handbook). In contrast, we assign relatively non-informative priors to $\bm{\beta}$, $\lambda$, and the neural network weight and bias parameters $\bm{\theta}_{W}$, with $\bm{\mu}_{\bm{\beta}}=\bm{\mu}_{W}=\bm{0}$, $c_{W}=\sigma^{2}_{\lambda}=1$, $\mu_{\lambda}=0$, and $c_{\bm{\beta}}=10$. The neural network for ${\bm{v}}_{NN,t}$ consists of $L=3$ hidden layers, each with 16 hidden units.
@@
As discussed in Section 4.2, the specification of the B-PINN model in this BHM uses “pseudo-data” for the PDE residuals, boundary conditions, and initial conditions. Specifically, we can write the following conditional data distributions that have the effect of updating the prior distributions for weights $\bm{\theta}_{W}$.
$[D_{r}|\sigma^{2}_{r},\bm{\theta}_{W},\lambda]$ $=\text{N}(0;[\frac{\partial v_{NN}}{\partial t}+v_{NN}\frac{\partial v_{NN}}{\partial s}-\lambda\frac{\partial^{2}v_{NN}}{{\partial s^{2}}}](s,t;\bm{\theta}_{W}),\sigma^{2}_{r}),\quad(\text{Burgers' equation})$
$[D_{bc+}|\sigma^{2}_{bc},\bm{\theta}_{W}]$ $=\text{N}(0;v_{NN}(\pi,t;\bm{\theta}_{W}),\sigma^{2}_{bc}),\quad(\text{Right boundary condition})$
$[D_{bc-}|\sigma^{2}_{bc},\bm{\theta}_{W}]$ $=\text{N}(0;v_{NN}(-\pi,t;\bm{\theta}_{W}),\sigma^{2}_{bc}),\quad(\text{Left boundary condition})$
$[D_{ic}|\sigma^{2}_{ic},\bm{\theta}_{W}]$ $=\text{N}(\exp(-s^{2});v_{NN}(s,0;\bm{\theta}_{W}),\sigma^{2}_{ic}),\quad(\text{Initial condition}),$
@@
where $\sigma^{2}_{r}$, $\sigma^{2}_{bc}$, and $\sigma^{2}_{ic}$ govern the extent to which the neural network $v_{NN}$ is constrained to satisfy Burgers’ equation along with the initial conditions and boundary conditions. Assuming reasonably strong prior information for the PDE, boundary conditions, and initial conditions, we set $\sigma^{2}_{r}=0.05^{2}$ and $\sigma^{2}_{bc}=\sigma^{2}_{ic}=0.01^{2}$. We re-emphasize that assuming the residual data are zero (e.g., $D_{r}=0$) further constrains the prior on $\bm{\theta}_{W}$ and these residuals are not actual measurements of the process.
Figure 3: Violin plots of the BHM model parameter posterior distributions, with the true value indicated by a dashed red line. From left to right, the plots correspond to $\sigma^{2}_{\delta}$, $\ell_{\delta}$, $\lambda$, $\sigma^{2}_{z}$, $\beta_{1}$, and $\beta_{2}$.
@@
After specifying the prior distributions and the neural network architecture, we draw 2,000 posterior samples using HMC with the No-U-Turn Sampler (10.5555/2627435.2638586), discarding the first 1,000 as burn-in and using a tree depth of 10. This is implemented using the “NumPyro" (phan2019composable) and “JAX" (jax2018github) packages in Python and takes approximately 45 minutes on an Apple MacBook M4 Pro with 128GB RAM.
@@ 6.1 Results
Figure 4: Left: true $\bm{v}_{t}$; Second from the left: predicted $v_{NN}(s,t)$; Second from the right: posterior means (dashed line) and corresponding 95% credible intervals (shaded region) for predicted values of a marginal spatial slice of $\bm{v}_{NN,t}$ at times $t=0$, $2.5$, and 55 compared to the true process (solid line); Right: Posterior means (dashed line) and corresponding 95% credible intervals (shaded region) for a marginal time slice of the predicted values of $\bm{v}_{t}$ at spatial locations s=3.14s=-3.14, 1.13-1.13, and $1.01$ compared to the true process (solid line). In the right two plots, the dots correspond to spatial locations or time points when data are observed.
Figure 5: First Row, First Column: true values of the spatio-temporal discrepancy term $\bm{\delta}_{t}$; Second Row, First Column: corresponding posterior prediction of $\bm{\delta}_{t}$; First Row, Columns 2-4: spatial slice of posterior mean (dashed line) and 95% credible intervals (shaded region) for times $t=0$, 22, and $t=5$ compared to the true process (solid line); Second Row, Columns 2-4: time slice of posterior mean (dashed line) and 95% credible intervals (shaded region) for spatial locations s=3s=-3, 1-1, and 1 compared to the true process (solid line). For all plots in columns 2-4, the dots correspond to spatial locations or time points when data is observed.
Figures 4 and 5 compare the true values and predictions of ${\bm{v}}_{t}$ and $\bm{\delta}_{t}$, respectively. As shown in Figure 4, the predicted ${\bm{v}}_{t}$ space-time field and spatial and temporal marginal plots align well with the true values, despite the substantial amount of missing data. This is one of the primary advantages of incorporating mechanistic information into the model. In contrast, Figure 5 demonstrates that accurately predicting the discrepancy term, $\bm{\delta}_{t}$, is more challenging and subject to greater uncertainty, although the posterior mean (dashed line in the spatial and temporal slice plots) still reasonably matches the true process (solid line). Combining posterior samples from each term, Figure 6 shows the posterior mean of the overall process $\bm{u}_{t}$. While there are slight deviations from the true $\bm{u}_{t}$, the posterior mean successfully captures the underlying spatio-temporal pattern, demonstrating the flexibility of the model.
We demonstrate a similar BHM-PINN for the case of normally-distributed data in the supplementary material. The results are quite comparable, if not slightly better, than those presented here for the Poisson data.
Figure 6: Top Left: true simulated $\exp(\bm{u}_{t})$; Top Right: predicted posterior mean of $\exp(\bm{u}_{t})$; Bottom Left: posterior standard deviation, with white rectangles indicating locations where data are observed; Bottom Right: difference between the true $\bm{u}_{t}$ and the posterior mean.
@@ 7 Discussion
The incorporation of mechanistic information in data-driven models has been of interest to scientists for decades. There are numerous paradigms for such data-driven approaches, ranging from variational data assimilation, online filtering, physical-statistical models and probabilistic numerical methods. Recently, there has been a resurgence of interest in these approaches because of the success of using deep neural networks with physical constraints (Raissi2020; yang2021b; zou2024correcting). Here, we review these PINN approaches and show some connections to the statistical approaches, followed by a generalization of B-PINNs to show that a general BHM framework can accommodate the B-PINN dynamical process but with the added flexibility to incorporate non-normal data models, covariate effects, and discrepancy processes. We illustrate such a BHM-PINN with the simulation of Poisson spatio-temporal data that are informed by the classic Burgers’ nonlinear PDE, a covariate-based mean process, and a GP discrepancy term, and demonstrate we can recover important components of the model given uncertain and missing data.
The main contribution of this work is to place the PINN and B-PINN in the context of more traditional statistical approaches to incorporate mechanistic information in spatio-temporal models. Importantly, as a generalization, we show that a B-PINN naturally fits into a BHM framework (i.e., a BHM-PINN), which can accommodate non-normal data, covariates, and random discrepancy terms.
@@
In some applications it maybe be preferable to embed a mechanistically-informed dynamical random spatio-temporal dynamical process as opposed to a neural network as in traditional physical-statistical BHMs. For example, wikle2001spatiotemporal model 2-D spatio-temporal wind velocity components as a vector-autoregressive process that is motivated by an underlying set of (linear) PDEs, and gopalan2019hierarchical motivates the latent process by a numerical PDE solution summed with a stochastic discrepancy process. However, for general nonlinear PDEs there may not be a convenient autoregressive or general quadratic nonlinear (wikle2010) representation of the latent dynamics. In such cases, a mechanistic-BHM using a B-PINN approach may be preferred since automatic differentiation can be used for general, nonlinear PDEs. This could also be an advantage of using a model such as the one in Section 5 over PNM methods that lack a neural network representation, in which case it is not obvious that automatic differentiation would apply.
Related to the previous point, in the classical DSTM BHM (e.g., cressie2011statistics) it is often useful to consider the latent dynamics in a dimension reduced framework such as ${\bm{v}}_{t}={\bm{\Phi}}{\bm{\alpha}}_{t}$ instead of neural network dynamics, ${\bm{v}}_{NN,t}$. In such a case, ${\bm{v}}_{t}$ is the $n$-dimensional latent dynamic process and $\bm{\Phi}$} is an $n\times n_{\alpha}$ dimensional basis function matrix with ${\bm{\alpha}}_{t}$ an $n_{\alpha}$-dimensional dynamic process where $n_{\alpha}<n$. For certain basis functions (e.g., Fourier) and systems, one can specify underlying mechanistic differential equations that govern the dynamics of ${\bm{\alpha}}_{t}$ (e.g., cressie2011statistics). This spectral approach could provide additional computational efficiency if the dynamics can be represented in a lower-dimensional manifold (reduced rank) basis expansion.
@@
From our exploration of these methods, in particular, the incorporation of a B-PINN within a more general BHM, there are several challenges with PINN-based models that suggest additional research. Of primary concern is the challenge of computational efficiency when scaling to a realistic spatio-temporal application. Although automatic differentiation is quite helpful for implementing PDE models, it can be slow, and there are well-known challenges to Bayesian learning of deep neural network parameters. In addition, real-world processes are often described by more complex dynamics than can be parameterized by a PDE (all models are wrong), and modeling the discrepancy becomes crucial. Our simulation studies have shown there can be significant identifiability issues with a flexible discrepancy process and the underlying mechanistic process without rough a priori guidance as to at least the space and time scales corresponding to each. It is also the case the data volume and orientation can make a difference in accurately estimating the various model components. Lastly, as with all deep neural implementations, there is typically no a priori guidance to inform the model architecture. These are very active areas of research and we are optimistic that these challenges can be met and these methods can be scaled to many realistic spatio-temporal problems.
@@ Acknowledgments
Supplementary Material
Figure 7: Simulated latent process $u(s,t)$ (left) and components over one-dimensional space and time: Burgers’ equation v(s,t){v}(s,t) (second from left), mean process $\mu(s,t)$ (second from right), and discrepancy process $\delta(s,t)$ (right).
We present a simulation experiment where the data model follows a Normal distribution as opposed to Poisson as in the main paper. Specifically, the details of the simulation are the same as those described in the manuscript, except for:
  1. 1.
  2. 2.
    Simulated Observations ($z(s,t)$): We then simulated observations from a Normal distribution, $z(s,t)\sim\text{N}(u(s,t),\sigma^{2}_{d}),$ where $\sigma^{2}_{d}=0.2$. We assumed that 50% of the space coordinates are missing.
(a) Violin plots of the BHM model parameter posterior distributions, with the true value indicated by a dashed red line. From left to right, the plots correspond to $\sigma^{2}_{\delta}$, $\ell_{\delta}$, $\lambda$, $\sigma^{2}_{d}$, $\beta_{1}$, and $\beta_{2}$.
(b) Left: true $\bm{v}(s,t)$; Second from the left: Posterior mean of $\bm{v}_{NN}(s,t)$; Second from the right: Posterior means (dashed line) and corresponding 95% credible intervals (shaded region) for predicted values of a marginal spatial slice of $\bm{v}_{NN}(s,t)$ at times $t=0$, $2.5$, and 55 compared to the true process (solid line); Right: Posterior means (dashed line) and corresponding 95% credible intervals (shaded region) for a marginal time slice of the predicted values of $\bm{v}_{NN}(s,t)$ at spatial locations s=3.14s=-3.14, 1.13-1.13, and $1.01$ compared to the true process (solid line). In the right two plots, the dots correspond to spatial locations or time points when data are observed.
Figure 9: Results
Figure 7 shows the simulated latent process and associated components and 8 shows the simulated data and the missingness used when fitting the model. Specifically, we consider the following BHM to fit these simulated data. First, using the notation defined in the main paper, the data and process models are given by
$\text{Data Model}:$ $\quad\bm{z}_{t}=\bm{H}_{t}\bm{u}_{t}+\bm{\epsilon}_{t},\quad\bm{\epsilon}_{t}\sim\text{N}(\bm{0},\sigma^{2}_{d}\bm{I})$
$\text{True Process}:$ $\quad u(s,t)=\bm{x}(s,t)^{\top}\bm{\beta}+v_{NN}(s,t)+\delta(s,t),$
where $\bm{H}_{t}$ is an incidence matrix. The latent dynamic and discrepancy processes are specified as
$\text{Latent dynamic process ($v_{NN}(s,t;\bm{\theta}_{W})$}:$ $\quad\frac{\partial{v}_{NN}(s,t)}{\partial t}\sim\text{N}(v_{NN}(s,t)\frac{\partial{v}_{NN}(s,t)}{\partial s}-\lambda\frac{\partial^{2}{v}_{NN}(s,t)}{\partial s^{2}},\sigma^{2}_{r})$
$\quad{v}_{NN}(s,0)\sim\text{N}(\exp(-s^{2}),\sigma^{2}_{ic})\quad\text{(Initial condition)}$
$\quad{v}_{NN}(-\pi,t),{v}_{NN}(\pi,t)\sim\text{N}(0,\sigma^{2}_{BC})\quad\text{(Boundary condition)}$
$\text{Discrepancy process}:$ $\quad\delta(s,t)\sim\text{GP}(0,k(s,s^{\prime}))$
$\quad k(s,s^{\prime})=\sigma^{2}_{\delta}\exp\bigg(-\frac{(s-s^{\prime})^{2}}{2\ell_{\delta}^{2}}\bigg).$
The prior distributions are specified as
$\sigma^{2}_{d}$ $\sim\text{Cauchy}(\mu_{d},\gamma_{d})\cdot\text{I}(\ell_{d}<\sigma^{2}_{d}<u_{d}),\qquad$ $\bm{\beta}\sim\text{N}(\bm{\mu}_{\bm{\beta}},c_{\bm{\beta}}\bm{I})$
$\lambda$ $\sim\text{N}(\mu_{\lambda},\sigma^{2}_{\lambda})\cdot\text{I}(0\leq\lambda<\infty),\qquad$ $\sigma^{2}_{\delta}\sim\text{Cauchy}(\mu_{\delta},\gamma_{\delta})\cdot\text{I}(\ell_{\delta}<\sigma^{2}_{\delta}<u_{\delta})$
$\ell_{\delta}$ $\sim\text{N}(\mu_{\ell},\sigma^{2}_{\ell})\cdot\text{I}(\ell_{\ell}\leq\ \ell_{\delta}<u_{\ell}),\qquad$ $\bm{\theta}_{W}\sim\text{N}(\bm{\mu}_{W},c_{W}\bm{I}),$
Figure 10: First Row, First Column: True values of the spatio-temporal discrepancy term $\bm{\delta}_{t}$; Second Row, First Column: Corresponding posterior prediction of $\bm{\delta}_{t}$; First Row, Columns 2-4: Spatial slice of posterior mean (dashed line) and 95% credible intervals (shaded region) for times $t=0$, 22, and $t=5$ compared to the true process (solid line); Second Row, Columns 2-4: Time slice of posterior mean (dashed line) and 95% credible intervals (shaded region) for spatial locations s=3s=-3, 1-1, and 1 compared to the true process (solid line). For all plots in columns 2-4, the dots correspond to spatial locations or time points when data is observed.
where we set $\mu_{d}=\mu_{\delta}=\mu_{\ell}=0,\gamma_{d}=\gamma_{\delta}=\sigma^{2}_{\ell}=1$, and truncation bounds $\ell_{d}=0.1,u_{d}=0.3,\ell_{\delta}=0.02,u_{\delta}=0.1,\ell_{\ell}=0.05,$ and $u_{\ell}=0.2$, in order to assign relatively informative priors for $\sigma^{2}_{d}$, $\sigma^{2}_{\delta}$, and $\ell_{\delta}$. For $\bm{\beta}$, $\lambda$, and neural network weight and bias parameters $\bm{\theta}_{W}$, we set $\bm{\mu}_{\bm{\beta}}=\bm{\mu}_{W}=\bm{0}$, $c_{W}=\sigma^{2}_{\lambda}=1$, $\mu_{\lambda}=0$, and $c_{\bm{\beta}}=10$. The neural network for ${\bm{u}}_{NN,t}$ consists of $L=3$ hidden layers, each with 16 hidden units. For PDE residuals, boundary, and initial condition data, we consider
$[D_{r}|\sigma^{2}_{r},\bm{\theta}_{W},\lambda]$ $=\text{N}(0;[\frac{\partial v_{NN}}{\partial t}+v_{NN}\frac{\partial v_{NN}}{\partial s}-\lambda\frac{\partial^{2}v_{NN}}{{\partial s^{2}}}](s,t;\bm{\theta}_{W}),\sigma^{2}_{r}),\quad(\text{Burgers' equation})$
$[D_{bc+}|\sigma^{2}_{bc},\bm{\theta}_{W}]$ $=\text{N}(0;v_{NN}(\pi,t;\bm{\theta}_{W}),\sigma^{2}_{bc}),\quad(\text{Right boundary condition})$
$[D_{bc-}|\sigma^{2}_{bc},\bm{\theta}_{W}]$ $=\text{N}(0;v_{NN}(-\pi,t;\bm{\theta}_{W}),\sigma^{2}_{bc}),\quad(\text{Left boundary condition})$
$[D_{ic}|\sigma^{2}_{ic},\bm{\theta}_{W}]$ $=\text{N}(\exp(-s^{2});v_{NN}(s,0;\bm{\theta}_{W}),\sigma^{2}_{ic}),\quad(\text{Initial condition}),$
where $\sigma^{2}_{r}=0.05^{2}$ and $\sigma^{2}_{bc}=\sigma^{2}_{ic}=0.01^{2}$.
After specifying the prior distributions and the neural network architecture, we draw 2,000 posterior samples using HMC with the No-U-Turn Sampler (10.5555/2627435.2638586), discarding the first 1,000 as burn-in and using a tree depth of 10. This is implemented using the “NumPyro" (phan2019composable) and “JAX" (jax2018github) packages in Python and takes approximately 45 minutes on an Apple MacBook M4 Pro with 128GB RAM.
Figure 9(a) presents violin plots for the parameters, excluding the neural network weights and biases $\bm{\theta}_{W}$, for which the credible intervals contain the true values. Figure 9(b) show comparisons between $v(s,t)$ and the posterior mean of $v_{NN}(s,t)$, showing that the posterior mean closely aligns with the true $v(s,t)$ despite a substantial amount of missing data. Figure 10 compares the true $\bm{\delta}_{t}$ with its predicted values, indicating that predicting the discrepancy term is more challenging subject to more uncertainty. Nevertheless, the posterior mean (dashed line in the spatial and temporal slice plots) provides a reasonably good match to the true process (solid line).
Figure 11 shows the posterior mean of the overall process $\bm{u}_{t}$. While there are minor deviations from the true $\bm{u}_{t}$, the posterior mean successfully captures the underlying spatio-temporal pattern, with the posterior standard deviation being higher in locations where data are missing. This demonstrates the flexibility of the model.